• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 17:35:11 on October 29, 1998

    Entry number 4055

    keyword=Beam Time Accounting
    Hall A

    10/29/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: A. Deur
    Shift Workers: A. Saha, B. Zhang

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1 36 1 1 55 1 1 40 7.18333333333333
    Beam Available
    but Not in Use
    0 24 0 0 0 0 0 20 0.733333333333333
    Beam Not Available
    or Unacceptable
    0 0 0 0 5 0 0 0 0.0833333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 36 1 1 1 1 1 40 7.26666666666667
    Planned Config-
    uration Change
    0 24 0 0 0 0 0 20 0.733333333333333
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 24 0 0 0 0 0 0 0.4
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 09:00 end last run @ E=3.4 GeV, Ep=1.875 GeV, theta=15.5. Tot number of events for |_ is 16.7 M 09:34 Bscope+newspot+harp scan. The last harp scan and the corrected BPMs value seem to be consistent. 9:50 begin runs // @ E=3.4 GeV, Ep=1.596 GeV, Theta=15.5 degree 15:15 NMR: target polarisation: 34 to 35 % 15:20. Change the target configuration to |_ 15: 35 begining of the |_ runs @E=3.4 GeV, p=1.596 GeV/c, Theta=15.5 degree

    Data Entry by: A. Deur